# Tricks from the Trade for Large-Scale Markdown Pricing: Heuristic Cut Generation for Lagrangian Decomposition

### v1

@@ 6 Experiments and Real-world Impact
Influence on Lagrangian term and original objective
It is important to note that the contributions to the objectives, in particular for the earlier iterations are dominated by the Lagrangian term $(b-Ax)^{\top}\lambda$ and not the primal objective $f(x)$ as shown in more detail in Fig. 3. Both the Lagrangian term and the primal objective improve substantially after the gap closes (in this example after the sixth iteration using the maximum violation heuristic in the solving strategy). [^2] Mathematically this means that the violation of feasibility and optimality are decreased at the same time. From a business perspective, this means that the solutions from later iterations are not only more aligned with business guidelines but also more profitable – highly desirable! At the same time the Lagrangian term remains much larger than the primal objective for the baseline solving strategy. Hence, it is plausible that for the baseline, when maximizing $\textnormal{LR}(\lambda^{j},X^{j})$, minimizing the violations of those constraints with large multipliers $\lambda^{j}$ will yield a larger dual than maximizing the primal objective.
Figure 3: Primal objective values (long term profit) of outer loop iterations vs value of the Lagrangian term in the dual across iterations. Primal objective value is the value found for the (relaxed) objective $f(x)$ in iteration $j$. Lagrangian term is the value of $(b-AX^{j})^{\top}\lambda^{j}$ of the dual solution in the iteration $j$. Compared are maximum violation heuristic and baseline.
We also observe in practice that we spend a lot of iterations exploring Lagrangian relaxations where at least some of the multipliers $\lambda^{j}_{\ell}$ are at the upper bound. The occurrence of the first $\lambda^{j}$ that does not contain any multipliers at $\bar{\lambda}$ coincides with the drop in the dual gap around iteration five (Fig. 2).
Finding solutions with smaller Lagrangian multipliers is desirable if such multipliers exist that still allow for minimal or no violation of relaxed constraints, as it shifts weights in the relaxed objective towards the primal objective. Large multipliers will create extreme solutions, which are undesirable in various ways. First they will often cause the other bound to be violated (as discussed earlier, we usually set upper and lower bound constraints). And second, at those extreme solutions all prices (and costs to the primal objective) are being paid in order to not violate those linking constraints with large Lagrangian multipliers, even if smaller multipliers would be sufficient. This is exactly what we see in Fig. 3: At the same time the Lagrangian term stops dominating the Lagrangian relaxation, the primal objective improves. For us, it is of particular importance to find solutions that have both small violations of the linking constraints and a large primal objective, since we intend to use those solutions in our primal heuristic given by Problem (5). In that case, finding solutions that are close to a dual optimum, and have high primal objectives is likely to directly translate into a better final solution (after the application of the primal heuristic). Again, we conclude that from our case study, the maximum violation heuristic seems to be the most promising heuristic cut-generation approach.
Trade-off between heuristics and iterations
We know that the maximum violation heuristic adds cuts in a fashion that is similar to the application of cutting planes to the Lagrangian dual (that is iterating between the solving of the Lagrangian relaxation given some multipliers and finding the multipliers that minimize the cutting plane problem). Comparing the effectiveness of the cuts by looking at the dual gap over the number of cuts added we see that heuristic-based cuts are slightly less effective in closing the gap than cuts that solve the Lagrangian relaxation; the baseline strategy converges 23 cuts earlier (at cut 49 vs 72). (Fig. 4).
This is consistent with the assumption that solving $\textnormal{LR}(\lambda)$ in each iteration provides the best bound and the best subgradient and hence adds the most efficient cut to the Cutting Plane problem. But practically, the difference in efficiency does not appear to be very large. The objective values and gradients supplied by combining previous solutions to the Lagrangian relaxation seem to provide good enough cuts to meaningfully update the cutting plane problem.
Figure 4: Relative dual gap of the cutting plane problem over time elapsed. Cuts include both those added based on the heuristic and from solving the LR. Each marker indicates the addition of a cut, and the text annotation show the number of cuts added so far. Baseline uses no heuristic to add cuts, heuristic uses the maximum violation heuristic 10 times per iteration (starting after the third iteration).
While the baseline strategy needs less iterations to converge, the time to convergence is sped up substantially when using our suggested strategy (Fig. 4). In our example, we observe a speed-up of more than 4.5 times (39 minutes vs 186 minutes). Looking across a larger set of experiments we see that the speedup is consistent in those tests (Table 1). Instead of solving the LR once, we can apply the maximum violation heuristic at least 20 times. The run-time requirement for the baseline solution are impractical in production, both because of time-constraints and due to computational costs.
avg time in seconds standard deviation
LR solving time 256.51 53.01
Time per heuristic 12.58 3.32
Table 1: Mean and standard deviation for solving LR or heuristic for one set of multipliers, computed on 23 independent test instances using 10 iterations for a total of 230 data points.
We conclude that our suggested strategy of using cuts generated via the maximum violation heuristic converges much faster to similarly good solutions than the baseline strategy. Thereby it mitigates the risk of being stuck with an non-converged solution when hitting the time limit imposed by business requirements (note how in Fig. 4 convergence takes place just around the critical three hour mark).
@@
For a fair comparison of the overall improvement that applying the maximum violation heuristic gives in a time-constrained setting, we went on to compare final solutions (after application of the primal heuristic described in Section 5) both with the maximum violation heuristic cut strategy (solving LR 10 times and heuristic 70 times) and without a heuristic cut strategy (LR 14 times).
Both strategies would take approximately equal time to solve. In this final experiment we want to investigate whether an improvement in solving the dual problem also leads to an improvement of the primal solution generated after applying our primal heuristic by solving problem (5).
@@
Table 2: Simulated improvement (in M€) of the primal solution in key KPIs for 8 start dates throughout the season. Improvements are the difference between the primal solution generated without the maximum violation heuristic, and the solution generated by applying the maximum violation heuristic 10 times per iteration. The last two rows show the average and the sum of the eight weeks
@@ 7 Conclusion
In this paper, we addressed the practical challenges encountered in implementing large-scale markdown strategies using a predict-then-optimize framework. These challenges primarily revolve around the stability and efficiency of identifying close-to-optimal solutions for huge mixed integer programs. We introduced novel heuristics that enhance the performance of existing methods, facilitating the rapid identification of high-quality solutions through a Lagrangian relaxation approach. An important component was the implementation of a heuristic cut generation scheme, based on a maximum-violation measure, and its seemless integration within an existing exact separation procedure. We demonstrated the practical impact of our methodology with empirical evidence from Zalando SE’s pricing systems. Our approach not only accelerates the process of finding high-quality solutions but also leads to multi-million revenue and profit increases, underscoring its commercial relevance.

### v2

@@ 6 Experiments and Real-world Impact
Comparing the maximum violation heuristic with partial aggregation formulations
As discussed earlier, solving the Lagrangian relaxation quickly is of utmost importance to us. Therefore, we wanted to compare directly the effectiveness of the maximum violation cut heuristic and compare it to the previously reported partial aggregation formulation, see Section 5 and  Guignard (2003); Frangioni (2005); Mitra et al. (2016). Importantly, we want to ensure that the maximum violation strategy produces as good or better bounds as the partial aggregation formulation would in similar time.
In order to compare these strategies, we first look at their performance in isolation. We take dual problems that were previously identified as ”difficult” as they required a large number of dual solutions (iterations) for the dual gap to close. For each of those problems we selected a set of dual solutions and applied one of two strategies:
  • For the maximum violation heuristic, we take the dual solutions and repeatedly apply the heuristic. Each repetition reduces the dual gap, at the cost of time. We can therefore determine the dual gap as a function of time.
  • For the partially aggregated problem, we solve the optimization problem for different group sizes, starting from a single group (creating the equivalent of baseline cutting plane problem) up to a fully disaggregated problem. The assignment of articles into groups is done randomly.
    Larger group sizes will lead to larger problems and thus slower solving times, but also usually result in tighter gaps. Hence, we can visualize the gap as a function of elapsed time.
This is equivalent to taking a set of solutions $X^{j}$ as described in Algorithm 1 and only applying steps 6 through 13. An example of a single problem is shown in Fig. 3. We can see that the maximum violation strategy closes the dual gap much quicker than the partially aggregated strategy for this example.
Figure 3: Example of applying the partial aggregation model or the maximum violation strategy to a fixed set of dual solutions. For maximum violation (red), the heuristic is applied repeatedly, and the observed dual gap $d_{j}$ and elapsed time are noted. For the disaggregated formulation (blue), different levels of aggregation are solved independently, and the elapsed solving time and dual gap is noted.
To test if this generalizes, we collected a set of 86 other ”difficult” dual problems. For all these instances, we were able to solve the fully disaggregated problem and thus obtained the best primal bound achievable given the parameters. As discussed in Section 5, the primal bounds provided by the disaggregated problem and by the maximum violation heuristic will eventually reach the same value. This is exactly what we see in our experiments: The bound provided by running the maximum violation heuristic until convergence was in all cases at the level of numeric precision of the best primal bound.
To compare the two strategies, we compute the time needed to reach a given gap to the best primal bound. For the heuristic, it is the time needed to achieve this gap by repeatedly applying the heuristic. For the partially aggregated strategy, we individually solve again the problem for different levels of aggregation and take the minimum of the solving times for the aggregations leading to a primal bound within the gap.
Looking at the geometric means across the tested samples (Tab. 1) we can see that the maximum violations strategy clearly outperforms the partially aggregated strategy in speed when solving to small gap sizes (0.1% and 0.01%), with the advantage increasing as the target gap gets smaller. For a gap of 0.01%, the geometric mean of solving times for the partially aggregated strategy is 48 s, while the maximum violation strategy takes less than 2 s. Interestingly, for a larger target gap of 1%, the geometric mean of solving times is nearly identical (0.12 s vs 0.11 s), but the geometric standard deviation is much larger for the partially aggregated problem.
gap to best bound partially aggregated maximum violation
0.1% 0.12 $\pm$ 18.43 0.11 $\pm$ 4.12
0.01% 4.27 $\pm$ 10.24 0.59 $\pm$ 3.45
0.001% 47.98 $\pm$ 3.35 1.69 $\pm$ 2.90
Table 1: Geometric mean $\pm$ geometric standard deviation of solving time (s) for different gaps to best primal bound across 86 examples.
The reason for this difference can be seen in the distribution of solving times for different gaps (Fig. 4). For larger target gaps (10% or 1%), the partially aggregated problem often solves them faster to that gap, with a lower median solve time, in cases where a partially aggregated problem with few groups was sufficient to solve to that target gap. At the same time, the longest solving times to the 10% and 1% gaps are much more extreme in the partially aggregated problems, sometimes exceeding 100 s, even though all problems took less than 20 s for the maximum violation heuristic. This split in the partially aggregated problems, where some problems are solved fast, while others are very slow, is also seen in the 0.1% gap target (with more problems taking longer than for the maximum violation heuristic). For the 0.01% target gap we see the heuristic very consistently outperforming.
Figure 4: Elapsed time to reach different gaps to best dual bound. For 86 different initial problems (a set of solutions to the relaxed dual problems, iterations from the Lagrangian decomposition) the partial aggregation and heuristic strategies were applied to find primal bounds. For partial aggregation, different group sizes were tested and used to determine the minimum time to gap. For the maximum violation heuristic, the heuristic was applied repeatedly, and the cumulative time to reach the target gap was recorded.
All-in-all, this leaves us with several clear advantages for the maximum violation heuristic over the partial aggregation strategy in our setup:
  • The maximum violation heuristic reaches tight primal bounds faster than the partial aggregation strategy. Particularly for tight bounds, the heuristic is more than 20 times faster (in geometric mean).
  • The maximum violation heuristic produces more consistent bounds at a given time. This consistency makes it easier to tune, as one can easier select how much time will usually result in ”tight enough” bounds.
  • Because the primal bound can be checked on every application of the maximum violation heuristic, it is easier to define stopping criteria.
Still, we want to compare how well these strategies perform when applying them to the complete solving of dual problems (that is, the full Algorithm 1). Here, we need to balance between the time spend obtaining dual solutions (solutions to $LR(\lambda^{j})$) and getting tight primal bounds and good corresponding Lagrangian multipliers $\lambda^{j}$. As we have seen in Fig. 4, this is particularly challenging for the partial aggregation strategy, as the quality of the bound for a given aggregation level can vary drastically between problems. Therefore, based on those results, we decided to try two different partial aggregation levels: the sub-problems are aggregated to either groups of 2000 (small) or 5000 (large) articles. Along with the maximum violation heuristic (adding 100 heuristic cuts on each multiplier evaluation) and the baseline (baseline cutting plane without either partial aggregation or cut heuristic), these strategies are applied to 3 problems that were previously found to be difficult to solve. The dual gaps for these problems are shown against the elapsed time in Fig. 5.
We see that the maximum violation heuristic (blue) provides a marked improvement over baseline (red), which fails to close the dual gap in all 3 examples. Interestingly, the maximum violation heuristic also outperforms the partially aggregated formulation for both aggregation levels. While both partially aggregated problems do provide reasonable gaps on the problems (with gaps at or below 0.01%), the maximum violation strategy reaches smaller gaps (around 0.0001%), and closes the gap faster. The maximum violation heuristic also takes less time per multiplier evaluation (outer loop in Algorithm 1) compared to the partially aggregated problems. Given that reaching very tight bounds consistently would require more than 100 s per multiplier evaluation, it is therefore unlikely that any further increasing of the group size would change the advantage we see for the maximum violation heuristic.
Figure 5: Full Lagrangian descent using different solving strategies. For 3 different problems (different weeks), the maximum violation strategy was applied (blue, 100 times per solving of the LR), and compared to baseline (red, baseline cutting plane), partially aggregated small (orange, creating 2000 random groups), and partially aggregated large (cyan, using 5000 groups). For each problem the dual gap $d_{j}$ was noted after each solving of the Lagrangian relaxation.
Based on these findings, i.e., quicker convergence, higher consistency and better tunability, we decided to select the maximum violation heuristic over the partial aggregation strategy for further experimentation and the real application.
@@
For a fair comparison of the overall improvement that applying the maximum violation heuristic gives in a time-constrained setting, we went on to compare final solutions (after application of the primal heuristic described in Section 5) both with the maximum violation heuristic cut strategy and without a heuristic cut strategy.
Since the use of the maximum violation heuristic cut strategy does not significantly increase run times, both strategies would take approximately equal time to solve and the results should be directly comparable. In this final experiment we want to investigate whether an improvement in solving the dual problem also leads to an improvement of the primal solution generated after applying our primal heuristic by solving problem (5).
@@
Table 2: Simulated improvement (in M€) of the primal solution in key KPIs for 8 start dates throughout the season. Improvements are the difference between the primal solution generated without the maximum violation heuristic, and the solution generated by applying the maximum violation heuristic 100 times per iteration. The last two rows show the average and the sum of the eight weeks
@@ 7 Conclusion
In this paper, we addressed the practical challenges encountered in implementing large-scale markdown strategies using a predict-then-optimize framework. These challenges primarily revolve around the stability and efficiency of identifying close-to-optimal solutions for huge mixed integer programs. We introduced novel heuristics that enhance the performance of existing methods, facilitating the rapid identification of high-quality solutions through a Lagrangian relaxation approach. An important component was the implementation of a heuristic cut generation scheme, based on a maximum-violation measure, and its seamless integration within an existing exact separation procedure. We demonstrated the practical impact of our methodology with empirical evidence from Zalando SE’s pricing systems. Our approach not only accelerates the process of finding high-quality solutions but also leads to multi-million revenue and profit increases, underscoring its commercial relevance.